OpenStudy (anonymous):

Fan and medal What is the solution to the system of equations? x=3y+5 x-4y=3 ( , )

OpenStudy (anonymous):

@arjund12

OpenStudy (anonymous):

You want me to teach you how to solve these types of question?

OpenStudy (anonymous):

put the 3y+5 in the first equation where you see x in the second equation and start with 3y+5−4y=3 then solve that one for y

OpenStudy (alexandervonhumboldt2):

as x=3y+5 and x=3+4y then 3y+5=3+4y

OpenStudy (alexandervonhumboldt2):

-y=-2 the y=2 now substitute y into one of the equations and solve for x

OpenStudy (anonymous):

3y+5−4y=3 take all the nos to one side n expressions to the other 3y−4y=3−5 −y=−2 therefore y=2 that you what y is solve for x x=3y+5 x=3(2)+5 and solve

OpenStudy (anonymous):

11

OpenStudy (anonymous):

2,11

OpenStudy (anonymous):

There's ur answer kiddo medal

OpenStudy (anonymous):

Thanks dude!

OpenStudy (anonymous):

np o.o